undeadknight26 (undeadknight26):

Israel started to solve a radical equation in this way: Square root of x plus 6 āˆ’ 4 = x Square root of x plus 6 āˆ’ 4 + 4 = x + 4 Square root of x plus 6 = x + 4 (Square root of x plus 6)2 = (x + 4)2 x + 6 = x2 + 8x + 16 x + 6 āˆ’ 6 = x2 + 8x + 16 āˆ’ 6 x = x2 + 8x + 10 x āˆ’ x = x2 + 8x + 10 āˆ’ x 0 = x2 + 7x + 10 0 = (x + 2)(x + 5)

undeadknight26 (undeadknight26):

Israel started to solve a radical equation in this way: Square root of x plus 6 āˆ’ 4 = x Square root of x plus 6 āˆ’ 4 + 4 = x + 4 Square root of x plus 6 = x + 4 (Square root of x plus 6)2 = (x + 4)2 x + 6 = x2 + 8x + 16 x + 6 āˆ’ 6 = x2 + 8x + 16 āˆ’ 6 x = x2 + 8x + 10 x āˆ’ x = x2 + 8x + 10 āˆ’ x 0 = x2 + 7x + 10 0 = (x + 2)(x + 5) x + 2 = 0 x + 5 = 0 x + 2 āˆ’ 2 = 0 āˆ’ 2 x + 5 āˆ’ 5 = 0 āˆ’ 5 x = āˆ’2 x = āˆ’5 Solutions = āˆ’2, āˆ’5 What error did Israel make? He subtracted 6 before subtracting x. He added 4 before squaring both sides. He factored x2 + 7x + 10 incorrectly. He did not check for extraneous solutions.

undeadknight26 (undeadknight26):

It is not B.

undeadknight26 (undeadknight26):

I think its A but not sure..I know he did C wrong but i just wanna see if he made a mistake in the early parts of it....I know he didn't get the right answer because the right answer is 2.

undeadknight26 (undeadknight26):

I know right?

OpenStudy (mathmath333):

He did not check for extraneous solutions.

OpenStudy (mathmath333):

Square root of x plus 6 āˆ’ 4 = x

OpenStudy (anonymous):

So, the answer is D, that he did not check for Extraneous solutions?